# RCP-Merging: Merging Long Chain-of-Thought Models with Domain-Specific Models by Considering Reasoning Capability as Prior

### v1

@@ Introduction
We demonstrate that RCP-Merging, requiring only a small number of open-source calibration samples, can effectively integrate long CoT reasoning capabilities into a domain-specific model. Through extensive experiments across various tasks and model architectures like Qwen2.5 (Qwen 2024) and Llama3.1 (Grattafiori et al. 2024), our method consistently produces merged models that not only preserve domain-specific expertise but also exhibit surprisingly long CoT capabilities when addressing domain-specific questions, ultimately elevating their performance in certain domains. Notably, the average performance of the merged model on eight datasets improves by 9.5% and 9.2% compared with the state-of-the-art method on BioMedicine and Finance domains, respectively. Moreover, though model merging aims to find a comprehensive model that compromises the performance of original models, our method improved performance by 4.5% and 0.7% on PubMedQA and MedQA datasets (Jin et al. 2019, 2020), respectively, and improved performance by 0.5% on ConvFinQA dataset (Cheng, Huang, and Wei 2024) compared to the original domain-specific models. To sum up, our contributions include:
@@
To address the limitations of magnitude-based approaches, activation-based methods leverage data-driven signals, such as model activations or gradients on a small calibration set, to obtain a more nuanced understanding of parameter importance (Springenberg et al. 2015; Shrikumar, Greenside, and Kundaje 2017; Sundararajan, Taly, and Yan 2017; Michel, Levy, and Neubig 2019; Maini et al. 2023; Wang et al. 2023; Liu et al. 2024).
@@
$S_{i,k}^{t}$ $\approx\left\|\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}^{i}}\cdot\theta_{t}^{i}\right\|_{d=d_{t}^{k}}$ (7)
$\approx\|g_{i,d_{k}}^{t}\cdot\theta_{t}^{i}\|_{d=d_{t}^{k}}.$
@@
$p^{i}_{r,t}=\left\|-\frac{1}{2N_{r}}\sum_{k=1}^{N_{r}}(g_{i,d_{k}}^{r})^{2}(\theta_{t}^{i}-\theta_{r}^{i})^{2}\right\|.$ (11)
@@
RCP-Merging achieves SOTA average performance on BioMedicine domain, surpassing all existing merging methods and even the original BioMedicine model meanwhile maintaining long CoT capability. The results, summarized in Table 1, demonstrate that RCP-Merging achieves a superior balance between domain-specific expertise and reasoning capabilities. It obtains the highest average score of 49.4 across all benchmarks, significantly surpassing all other merging methods. Specifically, in the target BioMedicine domain, RCP-Merging achieves top scores on both PubMedQA and MedQA with scores of 55.5 and 54.1, effectively integrating the knowledge from the BioMedicine model.
@@
Figure 3: Gibberish rate comparison of merging Qwen2.5-7B (Base), Meditron3-Qwen2.5-7B (BioMedicine) and DeepSeek-R1-Distill-Qwen-7B (Reasoning) on all datasets. The gibberish rate is the percentage of gibberish or nonsensical content generated for a given benchmark, and a lower value means higher-quality content.
@@
To address the emerging issue of model stability in merging methods, we measure the “gibberish rate”: the frequency of nonsensical outputs identified by a GPT-4 based evaluator (OpenAI 2023). This metric validates whether benchmark performance is genuine or from output degeneration.
As shown in Figure 3, RCP-Merging demonstrates superior stability, achieving the lowest average gibberish rate of 14.3%, with rates of 0% on PubMedQA and 5.8% on MedQA. In stark contrast, baseline methods like TIES and DARE TIES & AIM suffer from output collapse, showing high gibberish rates of 82.3% and 79.5% respectively. This confirms RCP-Merging’s robust performance stems from a genuine integration of capabilities. Moreover, we further evaluate the stability of RCP-Merging’s output using Distinct-N and PPL metrics, results are shown in Table 5 in Appendix.
@@
RCP-Merging demonstrates top performance across all evaluated categories, including GSM8k in Math, HumanEval in Code, ARC-C in Knowledge, and ConvFinQA in Finance (Cheng, Huang, and Wei 2024). Shown in Table 2, results demonstrate that RCP-Merging achieves the highest average score of 72.2, decisively outperforming all baseline methods on Finance domain. As the performance shown in the table, RCP-Merging demonstrates the best performance across four benchmarks, this further verifies the scalability of RCP-Merging in different fields. Results show RCP-Merging balances domain-specific performance and long CoT capability across multiple domains.
@@
We use GSM8k, HumanEval, ARC-C and PubMedQA to indicate the performance of different merge methods on Math, Code, Knowledge and BioMedicine domain. As the results in Table 3, RCP-Merging achieves the best average score of 68.3 among all merging techniques. Although FuseLLM with AIM shows a slightly better score in the specific BioMedicine domain, RCP-Merging has the best overall capability. We evaluate our method on different model size, with results shown in Table 7 in Appendix.
@@
Ablation Study
@@ B.3 Datasets and Evaluation Metrics
We conduct hyperparameter experiments for the reasoning-preserving coefficient $\lambda$ on the Qwen2.5-7B base architecture to investigate its impact on model merging performance. The primary goal is to identify an optimal $\lambda$ that effectively balances the integration of domain-specific knowledge with the preservation of foundational reasoning abilities. Our experimental results indicate that a $\lambda$ value of 0.3 yields the best overall performance across both medical and reasoning benchmarks for the 7B model.
The detailed results are illustrated in Figure 3. As shown in Figure 4(a), the merged model’s performance on the BioMedicine Benchmark peaks at $\lambda=0.3$, achieving accuracies of 55.5% on PubMedQA and 54.2% on MedQA. Notably, this represents a significant improvement over the specialized BioMedicine model on PubMedQA. For the Reasoning Benchmark, depicted in Figure 4(b), increasing $\lambda$ tends to degrade performance as it more heavily penalizes deviations from the reasoning model. However, at $\lambda=0.3$, the model maintains strong performance, achieving the highest accuracy for the merged model on both GSM8k and HumanEval. When $\lambda$ increases beyond 0.3, performance on all benchmarks declines, suggesting that an excessively high penalty compromises the model’s ability to effectively absorb new domain knowledge.

### v2

@@ Abstract
Codehttps://github.com/ZeroNLP/RCP-Merging
Datasetshttps://github.com/ZeroNLP/RCP-Merging
@@ Introduction
We demonstrate that RCP-Merging, requiring only a small number of open-source calibration samples, can effectively integrate long CoT reasoning capabilities into a domain-specific model. Through extensive experiments across various tasks and model architectures like Qwen2.5 (Qwen 2024) and Llama3.1 (Grattafiori et al. 2024), our method consistently produces merged models that not only preserve domain-specific expertise but also exhibit surprisingly long CoT capabilities when addressing domain-specific questions, ultimately elevating their performance in certain domains. Notably, the average performance of the merged model on eight datasets improves by 9.5% and 9.2% compared with the state-of-the-art method on BioMedicine and Finance domains, respectively. Moreover, though model merging aims to find a comprehensive model that compromises the performance of original models, our method improved performance by 4.5% and 0.7% on PubMedQA and MedQA datasets (Jin et al. 2019, 2020), respectively, and improved performance by 0.5% on ConvFinQA dataset (Cheng et al. 2024) compared to the original domain-specific models. To sum up, our contributions include:
@@
To address the limitations of magnitude-based approaches, activation-based methods leverage data-driven signals, such as model activations or gradients on a small calibration set, to obtain a more nuanced understanding of parameter importance (Springenberg et al. 2015; Shrikumar et al. 2017; Sundararajan et al. 2017; Michel et al. 2019; Maini et al. 2023; Wang et al. 2023; Liu et al. 2024).
@@
$S_{i,k}^{t}$ $\approx\left\|\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}^{i}}\cdot\theta_{t}^{i}\right\|_{d=d_{t}^{k}}\approx\|g_{i,d_{k}}^{t}\cdot\theta_{t}^{i}\|_{d=d_{t}^{k}}.$ (7)
@@
$p^{i}_{r,t}=-\left\|\frac{1}{2N_{r}}\sum_{k=1}^{N_{r}}(g_{i,d_{k}}^{r})^{2}(\theta_{t}^{i}-\theta_{r}^{i})^{2}\right\|.$ (11)
@@
RCP-Merging achieves SOTA average performance on BioMedicine domain, surpassing all existing merging methods and even the original BioMedicine model meanwhile maintaining long CoT capability. The results, summarized in Table 1, demonstrate that RCP-Merging achieves a superior balance between domain-specific expertise and reasoning capabilities. It obtains the highest average score of 49.4 across all benchmarks, significantly surpassing all other merging methods. Specifically, in the target BioMedicine domain, RCP-Merging achieves top scores on both PubMedQA and MedQA with scores of 55.5 and 54.1, effectively integrating BioMedicine knowledge.
@@
Figure 3: Gibberish rate comparison for merging Qwen2.5-7B (Base), Meditron3-Qwen2.5-7B (BioMedicine), and DeepSeek-R1-Distill-Qwen-7B (Reasoning) on all datasets, where a lower rate indicates higher-quality content.
@@
To address model stability, we measure gibberish rate: the frequency of nonsensical outputs identified by a GPT4 evaluator (OpenAI 2023) to validate genuine performance against output degeneration. As shown in Figure 3, RCP-Merging demonstrates superior stability, achieving a low 14.3% average gibberish rate (0% on PubMedQA, 5.8% on MedQA). This starkly contrasts with baseline methods like TIES (82.3%) and DARE TIES AIM (79.5%), which suffer from significant output collapse. This confirms RCP-Merging’s robust performance stems from genuine capability integration.
@@
RCP-Merging demonstrates top performance across all evaluated categories, including GSM8k in Math, HumanEval in Code, ARC-C in Knowledge, and ConvFinQA in Finance (Cheng et al. 2024). Shown in Table 2, results demonstrate that RCP-Merging achieves the highest average score of 72.2, decisively outperforming all baseline methods on Finance domain. As the performance shown in the table, RCP-Merging demonstrates the best performance across four benchmarks, this further verifies the scalability of RCP-Merging in different fields. Results show RCP-Merging balances domain-specific performance and long CoT capability across multiple domains.
@@
We use GSM8k, HumanEval, ARC-C and PubMedQA to indicate the performance of different merge methods on Math, Code, Knowledge and BioMedicine domain. As the results in Table 3, RCP-Merging achieves the best average score of 68.3 among all merging techniques. Although FuseLLM with AIM shows a slightly better score in the specific BioMedicine domain, RCP-Merging has the best overall capability.
@@
Hypterparameter Analysis.
(a) BioMedicine Benchmark
(b) Reasoning Benchmark
Ablation Study.
@@ Conclusion
Acknowledgement
This work was supported by National Natural Science Foundation of China (62406114,62472181,62306117), the Fundamental Research Funds for the Central Universities (2024ZYGXZR074), Guangdong Basic and Applied Basic Research Foundation (2025A1515011413,2024A04J3681,2024A1515010220), and GJYC program of Guangzhou (2024D03J0005), National Key R & D Project from Minister of Science and Technology (2024YFA1211500), and CCF-Baidu Open Fund.